OpenStudy (anonymous):

Find exact value of cos(A-B) given: tan A= 2/5 if angle A is in quadrant 1 and cos B=-2/3 if angle B is in quadrant 3.

OpenStudy (mertsj):

\[\cos (a-b)=\cos a \cos b+\sin a \sin b\]

OpenStudy (anonymous):